OpenStudy (timaashorty):

I'm stuck on a few calculations for Heat of combustion lab. Can anyone please help me?

OpenStudy (aaronq):

what question?

OpenStudy (timaashorty):

@aaronq

OpenStudy (aaronq):

hm can you post it as a pdf? i don't like downloading files

OpenStudy (timaashorty):

How can I do that?

OpenStudy (timaashorty):

how about a snapshot?

OpenStudy (aaronq):

yeah that works. as long as i don't have to download it

OpenStudy (timaashorty):

I'm not sure if you have to download it, but give me a sec. ( and I don't worry it's safe ;P)

OpenStudy (aaronq):

haha okay.

OpenStudy (timaashorty):

Yup ; no download (:

OpenStudy (timaashorty):

I need help with 2 in Calibration of calorimeter & 3, 4 in Heat combustion of cruscose.

OpenStudy (timaashorty):

*sucrose

OpenStudy (aaronq):

for #2, simply use the formula given \(\Delta H=C*\Delta T\) \(\Delta H\) is given as 3226 kJ/mol

OpenStudy (timaashorty):

so 3226 kJ/mol / triangle T ?

OpenStudy (aaronq):

yes

OpenStudy (aaronq):

oh wait i didn't see the n

OpenStudy (timaashorty):

and to just make sure, triangle t is the answer for number one correct?

OpenStudy (aaronq):

it's ΔH=C∗ΔT/n, so C=\(\dfrac{ΔH*n}{ΔT}\)

OpenStudy (timaashorty):

What does the n stand for?

OpenStudy (aaronq):

moles

OpenStudy (timaashorty):

wait, whats triangle h?

OpenStudy (aaronq):

change in Enthalpy

OpenStudy (timaashorty):

how do i find that?

OpenStudy (aaronq):

it's given in the question

OpenStudy (timaashorty):

so .. 3226 kJ/mol * 0.008787/ 2.887 = 9.819 kj/mol ?

OpenStudy (aaronq):

well mol is cancelling out, and you have /K so the units should be kJ/K

OpenStudy (timaashorty):

Oh okay, but the calculations is correct?

OpenStudy (aaronq):

yep

OpenStudy (timaashorty):

Thanks, can you show help with 3 ?

OpenStudy (aaronq):

#3 is pretty much the same thing, just different numbers

OpenStudy (aaronq):

and you're solving for \(\Delta H\)

OpenStudy (timaashorty):

answer for number 2 times trianlge t / moles? x 9.819 * 2.45/ 0.002888 ?

OpenStudy (timaashorty):

ignore the x.

OpenStudy (aaronq):

yes

OpenStudy (aaronq):

the last part is pretty self explanatory. do you know how to do it?

OpenStudy (timaashorty):

the answer I got divided by the correct answer x 100 = the percentage of error? 8329.830/ 5639 x 100 = 147.7 % D: wut.

OpenStudy (aaronq):

you're missing the subtraction (8329.830-5639)/5639*100

OpenStudy (timaashorty):

47.7 % ?

OpenStudy (aaronq):

yeah

OpenStudy (timaashorty):

Thank you so much! You're a really great helper (:!

OpenStudy (aaronq):

thanks! and no problem at all